### v1

# When isometry and equivalence for skew constacyclic codes coincide

Abstract.
We show that the notions of $(\mathbf{n},\sigma)$-isometry and $(\mathbf{n},\sigma)$-equivalence introduced by Ou-azzou et al coincide for most skew $(\sigma,a)$-constacyclic codes of length $\mathbf{n}$. To prove this, we show that all Hamming-weight-preserving homomorphisms between their ambient algebras must have degree one when those algebras are nonassociative. We work in the general setting of commutative base rings $S$. As a consequence, we propose new definitions of equivalence and isometry of skew constacyclic codes that exactly capture all Hamming-preserving isomorphisms, and lead to tighter classifications. In the process we determine homomorphisms between nonassociative Petit algebras, prioritizing the algebras $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$, which give rise to skew constacyclic codes.
Key words and phrases:
skew cyclic codes, skew polycyclic codes, quantum error-correcting codes, nonassociative algebra
Key words and phrases:
Skew polycyclic codes, skew polynomials, skew constacyclic codes, nonassociative algebras
2020 Mathematics Subject Classification:
Primary: 94B40, 94B05; Secondary: 17A35, 94B05
2020 Mathematics Subject Classification:
94B15, 11T71, 17A99, 81P70
The first author’s research is supported by NSERC Discovery Grant RGPIN-2025-05630.
1. Introduction
Petit algebras are nonassociative algebras that, when defined over finite fields or rings, are increasingly important for constructing and classifying skew polycyclic codes, as there is a one-one correspondence between skew polycyclic codes and principal left ideals in Petit algebras [Pum17]. Two such codes over a finite field are isometric if there is a Hamming-weight preserving isomorphism between them that comes from a monomial isomorphism between their corresponding Petit algebras viewed as algebras over the prime field, where the Hamming weight of a polynomial $p(t)$ is defined as the number of nonzero coefficients of $p(t)$. Such an isomorphism between Petit algebras, also called an isometry, thus gives a correspondence between their sets of associated codes that preserves code rate and Hamming distance as well as length and dimension [Pum25].
There has been significant work to understand these isometries. In [OaHA25, OaNA25], the authors consider a subclass of such isometries (namely, those monomial isomorphisms that act as the identity on the finite field $\mathbb{F}_{q}$) to propose the notions of $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry between skew constacyclic codes (and of $(\mathbf{n},\sigma)$-equivalence for skew polycyclic codes) over $\mathbb{F}_{q}$, generalizing [CDL14]. The same subclass of isometries was used, but between two associative algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$, to define equivalences of skew constacyclic codes in [BBB21] and later again in [LMn25]. Skew constacyclic codes are the skew polycyclic codes obtained from Petit algebras of the form $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ (notation as in Section 2) for $S$ a unital commutative (associative) ring and $a\in S^{\times}$.
In this paper we prove that these notions coincide for all classes of skew constacyclic codes over $S$ arising from proper nonassociative Petit algebras, a result that was only previously known for when the order $\mathbf{m}$ of $\sigma$ satisfied $\mathbf{m}\geq\mathbf{n}-1$ and when $S=K$ was a field [BP18].
The motivation for the present work was two-fold: firstly, we wanted to understand what turned out to be an error in [OaNA25, Theorem 4], which suggested the existence of isometries of higher degree for all skew constacyclic codes. Searching for these led us to discover a rather startling result.
To state it succinctly, note that we say that the algebra $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ is proper nonassociative when it is not an associative algebra; by [Pet67, (15)] this occurs exactly when $\sigma(a)\neq a$ or $\mathbf{m}\nmid\mathbf{n}$.
Theorem (Corollary 4.5).
Let $K$ be a field with $\sigma\in\operatorname{Aut}(K)$ of finite order $\mathbf{m}$, and let $a_{1},a_{2}\in K^{\times}$. If the Petit algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a_{i})$ are proper nonassociative, then all nonzero Hamming weight preserving homomorphisms $G$ between them such that $G|_{K}$ commutes with $\sigma$ are monomial of degree one, that is, satisfy $G(t)=\alpha t$ for some $\alpha\in K^{\times}$.
In particular, the theorem implies that these algebras admit no nonzero Hamming weight preserving homomorphisms of the form $G(t)=\alpha t^{k}$ for any $\alpha\in K^{\times}$ and $1<k<\mathbf{n}$.
We attack this question with a novel approach, and in an important greater generality, by first characterizing the power-associative monomials $\alpha t^{k}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, where now we allow $S$ to be any commutative unital (associative) ring, and $\alpha,b\in S^{\times}$. The answer is quite elegant.
Theorem 3.3.
The monomial $\alpha t^{k}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, with $\alpha\in S$ not a zero divisor and $1\leq k<\mathbf{n}$, is power-associative if and only if
(1.1) $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b).$
if we assume that $(\alpha t^{k})^{r}(\alpha t^{k})=(\alpha t^{k})(\alpha t^{k})^{r}$ for the least $r>0$ such that $rk\geq{\mathbf{n}}$.
The proof turns on a careful combinatorial argument which occupies most of Section 2. As a corollary, we give a classification of all monomial homomorphisms $G:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ for which the leading coefficient of $G(t)$ is not a zero divisor and for which the restriction of $G$ to $S$ is an automorphism commuting with $\sigma$ (Corollary 3.4).
We prove in Corollary 3.4 that if $k>1$ and $\tau$ is an automorphism of $S$ commuting with the order-$\mathbf{m}$ automorphism $\sigma$, then being a homomorphism $G:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ given by $G(a)=\tau(a)$ and $G(t)=\alpha t^{k}$ entails $\mathbf{m}|\mathbf{n}$, $k\equiv 1\mod\mathbf{m}$ and thus $\sigma(b)=b$, whence $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is associative.
We then leverage this to prove in Corollary 4.5 that if $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is proper nonassociative, then in fact there are no nonzero monomial homomorphisms $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ of degree other than one. Thus $(\mathbf{n},\sigma)$-isometry and $(\mathbf{n},\sigma)$-equivalence as introduced in [OaNA25] coincide for classes of skew constacyclic codes that are associated with proper nonassociative Petit algebras.
We then explain how our results lead to the failure of [OaNA25, Theorem 4] in Section 5, providing explicit counterexamples. We also recalculate the number of $(\mathbf{n},\sigma)$-equivalence classes of $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$, for $K$ a finite field, in Theorem 5.6.
Moreover, in Section 5 we go one to define more general notions of equivalence and isometry that perfectly capture the set of isomorphisms that will preserve Hamming distance. This offers a tighter classification of equivalence and isometry classes of codes (see Definition 5.1). The advantage of this perspective is clearly demonstrated in Example 5.8). For associative algebras, a variation of [OaNA25, Theorem 4] can be proved (Theorem 5.9). In particular, the existence of a monomial isomorphism of degree $k>1$ in the associative case gives rise to monomial isomorphisms of degree one with another algebra.
Section 6 addresses our second motivation for this work which is to eventually extend our parametrization of proper nonassociative Petit division algebras of small dimension that covered the case that $\mathbf{n}=\mathbf{m}$ is prime and $f(t)=t^{\mathbf{n}}-a$ [NP25]. To extend these results to any $\mathbf{n}$ and $\mathbf{m}$ we need to understand all possible isomorphisms between two division algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$, not only the monomial ones.
These division algebras are in one one correspondence with maximum rank distance codes. Any parametrization will thus yield parametrizations of the corresponding codes. For $\mathbf{n}=\mathbf{m}$, the division algebras $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ have just been used in a variation of Learning with Errors, a lattice based approach to postquantum cryptography [ML25].
When $S$ is an integral domain, we prove that polynomial, i.e., non monomial, morphisms between properly nonassociative algebras of the form $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ do not occur, subject to a technical hypothesis we denote ($\star$6.3).
As a by product, we obtain from Theorem 6.1 that for integral domains we obtain that if $a,b\in S^{\times}$, and $\mathbf{m}\nmid\mathbf{n}$, then any nonzero homomorphism $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$, must be monomial (Corollary 6.2).
Skew constacyclic and skew polycyclic codes over finite rings and mixed alphabets are currently among the most studied kinds of linear codes, for their applications to constructions of quantum error-correcting codes (such as via Gray maps and the CSS construction [BDU+20, VPSI23]), with many papers appearing on this topic in recent years, including for example [BP25, PIP22, YSP23, PPI25, RSG24]. Our current work serves to improve the de-duplication of codes by classifying all isometries between them, which will in turn make the search for good quantum error-correcting codes more efficient.
In particular, when considering the CSS construction for quantum error-correcting codes, one needs to concentrate on codes $C$ such that the dual code $C^{\perp}$ satisfies $C^{\perp}\subset C$ as explained in [DBU+21]. Thus an important further extension of our work is to consider when our isometries preserve some suitable notion of duality for these codes.
2. Petit algebras and power associativity
Let $S$ be a commutative, associative unital ring and suppose $\sigma\in\operatorname{Aut}(S)$ has finite order $\mathbf{m}$. Important examples in applications to coding theory are $S=K$ a finite field, or $S$ a chain ring (also called a finite local ring) or more generally a Galois ring (see, for example, [Pum17, §5.1]). In both these examples, every non-invertible element is a zero divisor.
Write $S_{0}=\{s\in S:\sigma(s)=s\}$ for the subring fixed by $\sigma$. If $\tau\in\operatorname{Aut}(S)$ is an arbitrary automorphism then we write $S_{0}^{\tau}=\{s\in S_{0}:\tau(s)=s\}$ for the common fixed subring under $\tau$ and $\sigma$. We define $K_{0},K_{0}^{\tau}$ analogously.
The ring $R=S[t;\sigma]$ is a noncommutative polynomial ring, with addition the usual addition of polynomials, and with multiplication defined by the relation
$t^{j}a=\sigma^{j}(a)t^{j}$
for all $a\in S$ and all $j\in\mathbb{N}$. It is an algebra over $S_{0}$.
Let $f(t)=t^{\mathbf{n}}-f_{0}(t)\in S[t;\sigma]$ be a monic polynomial of degree $\mathbf{n}\geq 1$. For all $g(t)\in S[t;\sigma]$ of degree $\ell\geq\mathbf{n}$, there exist uniquely determined $r(t),q(t)\in S[t;\sigma]$ with ${\rm deg}(r)<{\rm deg}(f)$, such that $g(t)=q(t)f(t)+r(t)$ [Pum17, Proposition 1]. Let $r:=g\,{\rm mod}_{r}f$ denote the remainder of this right division of $g$ by $f$. Since the remainders are uniquely determined, we may identify
$S[t;\sigma]/S[t;\sigma]f(t)=\{\sum_{i=0}^{\mathbf{n}-1}b_{i}t^{i}:b_{i}\in S\}$
as left $S[t;\sigma]$-modules. In fact, this quotient also carries the structure of a (potentially) nonassociative unital ring that is an algebra over the fixed ring $S_{0}$, with multiplication of any two polynomials of degree less than $\mathbf{n}$ defined by
$g\circ h=gh\,\,{\rm mod}_{r}f.$
In particular, for all $0\leq i,j<\mathbf{n}$ we have
$t^{i}t^{j}=\begin{cases}t^{i+j}&\text{if $i+j<\mathbf{n}$;}\\ t^{i+j-\mathbf{n}}f_{0}(t)&\text{if $i+j>\mathbf{n}$.}\end{cases}$
Note that $S[t;\sigma]/S[t;\sigma]f(t)$ is associative exactly when $S[t;\sigma]f(t)$ is a two-sided ideal. For example, this happens when $S=K$, $f(t)=t^{\mathbf{n}}-\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\in S_{0}[t]$ and $\sigma^{\mathbf{n}-i}=\operatorname{id}$ for all those $i$ where $a_{i}\not=0$ [Pet67, (15)].
We are particularly interested in the special case that $f(t)=t^{\mathbf{n}}-a$ for some $a\in S^{\times}$. Then the lack of power associativity is best understood by observing that in $S[t;\sigma]$, we have the identity $t^{\mathbf{n}+1}=t^{\mathbf{n}}t=tt^{\mathbf{n}}$ but in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$,
$t^{\mathbf{n}}t=at\quad\text{whereas}\quad tt^{\mathbf{n}}=ta=\sigma(a)t.$
That is, one must take care to simplify products only with right division by $t^{\mathbf{n}}-a$, as then it is clear that the correct reduction is
$t^{\mathbf{n}+1}\equiv t(t^{\mathbf{n}}-a+a)\equiv ta\equiv\sigma(a)t\text{ in }S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a).$
From this point onwards we write simply equality rather than congruence.
We require some additional notation. For any $i\in\mathbb{N}$, $\tau\in\operatorname{Aut}(S)$ and $\beta\in S$, write
$N_{i}^{\tau}(\beta)=\prod_{j=0}^{i-1}\tau^{j}(\beta).$
Note that $\beta\tau(N_{i}^{\tau}(\beta))=N_{i+1}^{\tau}(\beta)$ and thus
(2.1) $N_{i+j}^{\tau}(\beta)=N_{i}^{\tau}(\beta)\cdot\sigma^{i}(N_{j}^{\tau}(\beta))$
for all $i,j\geq 0$. If $S=K$ is a field, $K/K_{0}$ a cyclic Galois extension, and $\operatorname{Gal}(K/K_{0})=\langle\sigma\rangle$, where $\sigma$ has order $\mathbf{m}$, then $N_{\mathbf{m}}^{\sigma}$ is simply the norm map $N_{K/K_{0}}$ corresponding to $K/K_{0}$. More generally, $N_{\mathbf{m}}^{\sigma}(\beta)\in S_{0}$ for all $\beta\in S$ so we distinguish this case with the notation $N_{\mathbf{m}}^{\sigma}=N_{S/S_{0}}$.
Let us define the class of homomorphisms of Petit algebras that correspond to Hamming-weight preserving homomorphisms of skew polycyclic codes.
Definition 2.1.
Let $f,h\in S[t;\sigma]$ be monic polynomials of degree $\mathbf{n}$. Suppose $\tau\in\operatorname{Aut}(S)$, $\alpha\in S$ is nonzero and $k\in\mathbb{N}$. For $R=S[t;\sigma]$ or $R=S[t;\sigma]/S[t;\sigma]f(t)$, if there exists a ring homomorphism
$G_{\tau,\alpha,k}:R\to S[t;\sigma]/S[t;\sigma]h(t)$
defined via $G_{\tau,\alpha,k}|_{S}=\tau$ and $G_{\tau,\alpha,k}(t)=\alpha t^{k}$, that is,
(2.2) $G_{\tau,\alpha,k}\left(\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\right)=\sum_{i=0}^{\mathbf{n}-1}\tau(a_{i})(\alpha t^{k})^{i}.$
then we call it a monomial homomorphism of degree $k$. When $k=1$, we write $G_{\tau,\alpha}$ in place of $G_{\tau,\alpha,1}$, in which case we have simply
$G_{\tau,\alpha}\left(\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\right)=\sum_{i=0}^{\mathbf{n}-1}\tau(a_{i})N_{i}^{\sigma}(\alpha)t^{i}.$
As we shall discuss in Section 2, there is significant subtlety in the well-definedness of the term $(\alpha t^{k})^{i}$ in the nonassociative case.
In this paper, we expand the set of cases for which it is known that all homomorphisms or isomorphisms of Petit algebras are monomial. This was previously known in several cases, but so far only for isomorphisms.
Then the following results were originally stated for $S_{0}$-isomorphisms for the case that $S$ is a field in [BP18, Theorems 28, 29], and generalized the present setting in [Pum25]. To state them, set
$f(t)=t^{\mathbf{n}}-\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i},\quad h(t)=t^{\mathbf{n}}-\sum_{i=0}^{\mathbf{n}-1}b_{i}t^{i}.$
Theorem 2.2.
Let $\alpha\in S^{\times}$ and assume that $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. Then
$G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma]h(t)$
is an isomorphism of $S_{0}^{\tau}$-algebras if and only if
(2.3) $\tau(a_{i})=N_{\mathbf{n}-i}^{\sigma}(\sigma^{i}(\alpha))b_{i}$
for all $i\in\{0,\ldots,\mathbf{n}-1\}$.
When $K[t;\sigma]/K[t;\sigma]f(t)$ and $K[t;\sigma]/K[t;\sigma]h(t)$ are proper nonassociative, and $\operatorname{Aut}(K)$ is abelian, we can say even more.
Theorem 2.3.
Let $\mathbf{m}\geq\mathbf{n}-1$ and assume that $f$ and $h$ do not generate two-sided ideals in $K[t;\sigma]$. Suppose $\operatorname{Aut}(K)$ is abelian. For any subfield $F\subset K_{0}$, the only $F$-algebra isomorphisms
$G:K[t;\sigma]/K[t;\sigma]f(t)\to K[t;\sigma]/K[t;\sigma]h(t)$
are of the form $G=G_{\tau,\alpha}$ such that $\tau\in\operatorname{Aut}(K)$, $K_{0}^{\tau}\subset F$, and $\alpha\in K^{\times}$ satisfying Theorem 2.2.
In the following, while our results are valid for all $\mathbf{m},\mathbf{n}$, their novelty is in the open case that $\mathbf{m}<\mathbf{n}-1$, and that we do not require that $S$ is a field.
3. Monomial homomorphisms from $S[t;\sigma]$ to a Petit algebra
Homomorphisms between skew polynomial rings with $S$ a division algebra were studied in [LL92]. Surjective homomorphism between $S[t;\sigma]$ for general rings $S$ were considered in [Rim78].
Let $S$ be a commutative unital ring. We begin with general restriction on monomial homomorphisms $G:S[t;\sigma]\to S[t;\sigma]$.
Lemma 3.1.
Let $\tau\in\operatorname{Aut}(S)$, $k\in\mathbb{N}$, and $\alpha\in S$ such that $\alpha$ is not a zero divisor. Then the map generated by $G(c)=\tau(c)$ for all $c\in S$ and $G(t)=\alpha t^{k}$ defines a homomorphism of rings from $S[t;\sigma]$ to itself if any only if $\tau\sigma\tau^{-1}=\sigma^{k}$. In particular, if $\tau$ and $\sigma$ commute then this occurs if and only if $k\equiv 1\mod\mathbf{m}$.
Proof.
The map $G:S[t;\sigma]\to S[t;\sigma]$ generated by $G(c)=\tau(c)$ for all $c\in S$ and $G(t)=\alpha t^{k}$ is given by
$G(\sum_{i=0}^{\ell}a_{i}t^{i})=\sum_{i=0}^{\ell}\tau(a_{i})N_{i}^{\sigma}(\alpha)t^{ik},$
for any $\ell\in\mathbb{N}$ and $a_{i}\in S$. This map is $S_{0}^{\tau}$-linear and will be multiplicative if and only if $G(t)G(c)=G(\sigma(c))G(t)$ for all $c\in S$. We compute
$G(t)G(c)=\alpha t^{k}\tau(c)=\alpha\sigma^{k}(\tau(c))t^{k}$
and
$G(\sigma(c))G(t)=\tau(\sigma(c))\alpha t^{k}.$
The cancellability of $\alpha$ implies that this holds for all $c\in S$ if and only if $\tau\sigma=\sigma^{k}\tau$. In particular, if $\sigma$ and $\tau$ commute, we must have $\sigma=\sigma^{k}$, or $k\equiv 1\mod\mathbf{m}$. ∎
A special case of this result has appeared as early as [Rim78, Theorem 3].
From now on, we assume that $\tau\in\operatorname{Aut}(S)$ commutes with $\sigma$, and that $\alpha\in S^{\times}$. In this setting, the resulting homomorphism $G$ is an isomorphism if and only if $k=1$ and $\alpha\in S^{\times}$. In that case, for any $a\in S^{\times}$ and $\mathbf{n}\in\mathbb{N}$ we have
$G(t^{\mathbf{n}}-a)=N^{\sigma}_{\mathbf{n}}(\alpha)t^{\mathbf{n}}-\tau(a)=N^{\sigma}_{\mathbf{n}}(\alpha)(t^{\mathbf{n}}-N^{\sigma}_{\mathbf{n}}(\alpha^{-1})\tau(a)).$
Thus $G$ induces an isomorphism of $S_{0}^{\tau}$-algebras
$G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)\quad\text{where $\tau(a)=bN^{\sigma}_{\mathbf{n}}(\alpha)$.}$
In this section we show that these are all the Hamming-weight preserving homomorphisms from $S[t;\sigma]$ to $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$.
The first key obstruction for a map into $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ to be a homomorphism is the power-associativity of the term $\alpha t^{k}$. We require a technical lemma that will in Theorem 3.3 be used to reduce the question of associativity of powers of $\alpha t^{k}$ in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ to a straightforward relation between $\alpha$ and $b$. Write $\lfloor x\rfloor$ for the greater integer that is less than or equal to $x$.
Lemma 3.2.
Suppose $\mathbf{n},k\in\mathbb{N}$ and $\alpha,b\in S$ are not zero divisors. Suppose
(3.1) $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b).$
Then for all $s,\ell\in\mathbb{N}$ we have
(3.2) $\sigma^{[sk]_{\mathbf{n}}}(N_{\ell}^{\sigma^{k}}(\alpha))\prod_{j=1}^{z}\sigma^{sk-j{\mathbf{n}}}(b)=\sigma^{sk}(N_{\ell}^{\sigma^{k}}(\alpha))\prod_{j=1}^{z}\sigma^{sk-j{\mathbf{n}}}(\sigma^{\ell k}(b)),$
where $z=\lfloor\frac{sk}{{\mathbf{n}}}\rfloor$ and $sk]_{\mathbf{n}}=sk-z{\mathbf{n}}$ is the residue mod $\mathbf{n}$} of $sk$.
Proof.
Since $\alpha,b$ are not zero divisors the same is true of every image under a power of $\sigma$, and we may work in the total ring of fractions of $S$ obtained by localizing on the set of non zero divisors. Then the hypothesis (3.1) can be rephrased as
$\frac{\alpha}{\sigma^{\mathbf{n}}(\alpha)}=\frac{\sigma^{k}(b)}{b}.$
It follows that for any positive integer $z$ we have
$\frac{\alpha}{\sigma^{z{\mathbf{n}}}(\alpha)}=\prod_{i=0}^{z-1}\sigma^{i{\mathbf{n}}}\left(\frac{\alpha}{\sigma^{\mathbf{n}}(\alpha)}\right)=\prod_{i=0}^{z-1}\sigma^{i{\mathbf{n}}}\left(\frac{\sigma^{k}(b)}{b}\right)=\prod_{j=1}^{z}\sigma^{z{\mathbf{n}}-j{\mathbf{n}}}\left(\frac{\sigma^{k}(b)}{b}\right).$
Now let $\ell\in\mathbb{N}$. Applying the function $N_{\ell}^{\sigma^{k}}$ to both sides and then simplifying yields
$\prod_{i=0}^{\ell-1}\sigma^{ik}\left(\prod_{j=1}^{z}\sigma^{z{\mathbf{n}}-j{\mathbf{n}}}\left(\frac{\sigma^{k}(b)}{b}\right)\right)=\prod_{j=1}^{z}\sigma^{z{\mathbf{n}}-j{\mathbf{n}}}\left(\prod_{i=0}^{\ell-1}\sigma^{ik}\left(\frac{\sigma^{k}(b)}{b}\right)\right)=\prod_{j=1}^{z}\sigma^{z{\mathbf{n}}-j{\mathbf{n}}}\left(\frac{\sigma^{\ell k}(b)}{b}\right).$
That is, (3.1) implies that for all $z,\ell\in\mathbb{N}$ we have
$\frac{N_{\ell}^{\sigma^{k}}(\alpha)}{\sigma^{z{\mathbf{n}}}(N_{\ell}^{\sigma^{k}}(\alpha))}=\prod_{j=1}^{z}\sigma^{z{\mathbf{n}}-j{\mathbf{n}}}\left(\frac{\sigma^{\ell k}(b)}{b}\right).$
Now let $s\in\mathbb{N}$ and define $z$ by $sk]_{\mathbf{n}}=sk-z{\mathbf{n}}$. Applying $\sigma^{[sk]_{\mathbf{n}}}=\sigma^{sk-z{\mathbf{n}}}$ to both sides yields
$\frac{\sigma^{[sk]_{\mathbf{n}}}(N_{\ell}^{\sigma^{k}}(\alpha))}{\sigma^{sk}(N_{\ell}^{\sigma^{k}}(\alpha))}=\prod_{j=1}^{z}\sigma^{sk-j{\mathbf{n}}}\left(\frac{\sigma^{\ell k}(b)}{b}\right).$
Finally, cross-multiplying and simplifying yields the desired equality
$\sigma^{[sk]_{\mathbf{n}}}(N_{\ell}^{\sigma^{k}}(\alpha))\prod_{j=1}^{z}\sigma^{sk-j{\mathbf{n}}}(b)=\sigma^{sk}(N_{\ell}^{\sigma^{k}}(\alpha))\prod_{j=1}^{z}\sigma^{sk-j{\mathbf{n}}+\ell k}(b).$
We now demonstrate that the condition (3.1) is exactly the key to the power associativity of a monomial in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$.
Theorem 3.3.
Let $\alpha,b\in S$ and $\mathbf{n}$$\in\mathbb{N}$. Suppose neither $\alpha$ nor $b$ are zero-divisors and $1\leq k<{\mathbf{n}}$. Consider the magma generated by the monomial
$\mathsf{z}=\alpha t^{k}$
in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$. If $\mathsf{z}^{r}\mathsf{z}=\mathsf{z}\mathsf{z}^{r}$ for the least $r>0$ such that $rk\geq{\mathbf{n}}$, then
(3.3) $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b).$
Conversely, if (3.3) holds, then for every $s\in\mathbb{N}$, the expression $(\alpha t^{k})^{s}$ is well-defined in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ (that is, $\mathsf{z}$ is power associative) and we have
(3.4) $\mathsf{z}^{s}=N_{s}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{sk}{{\mathbf{n}}}\rfloor}\sigma^{sk-i{\mathbf{n}}}(b)t^{[sk]_{\mathbf{n}}}$
where $\lfloor\frac{sk}{\mathbf{n}}\rfloor=j$ when $j{\mathbf{n}}\leq sk<(j+1){\mathbf{n}}$ and $sk]_{\mathbf{n}}=sk-\lfloor\frac{sk}{\mathbf{n}}\rfloor\mathbf{n}$ denotes the residue of $sk$ mod $\mathbf{n}$ in the interval $0,\mathbf{n}-1]$.
Proof.
We first prove that if (3.3) holds, then $\mathsf{z}$ generates a monoid (that is, all powers are well-defined). Write $L(\mathsf{z},s)$ for the left-nested product
$\mathsf{z}(\mathsf{z}(\mathsf{z}(\cdots(\mathsf{z}))))$
of $\mathsf{z}$ with itself $s$ times. Our first step is to show that this is equal to (3.4) by induction on $s$. The case $s=1$ is given. Assume for some $s\geq 1$ that $L(\mathsf{z},s)$ is given by (3.4). We wish to show that
(3.5) $L(\mathsf{z},s+1)=N_{s+1}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{(s+1)k}{\mathbf{n}}\rfloor}\sigma^{(s+1)k-i\mathbf{n}}(b)t^{[(s+1)k]_{\mathbf{n}}}.$
We compute
$L(\mathsf{z},s+1)$ $=\mathsf{z}\cdot L(\mathsf{z},s)$
$=\alpha t^{k}\cdot N_{s}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-i\mathbf{n}}(b)t^{[sk]_{\mathbf{n}}}$
$=\alpha\cdot\sigma^{k}\left(N_{s}^{\sigma^{k}}(\alpha)\right)\sigma^{k}\left(\prod_{i=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-i\mathbf{n}}(b)\right)\cdot t^{k}t^{[sk]_{\mathbf{n}}}$
$=N_{s+1}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{(s+1)k-i\mathbf{n}}(b)\cdot t^{k}t^{[sk]_{\mathbf{n}}}.$
Note that since both $k$ and $sk]_{\mathbf{n}}$ are less than $\mathbf{n}$, we simply have
[…]
In the former case we are done. In the latter case, we apply the relation
$t^{k+[sk]_{\mathbf{n}}}=t^{k+[sk]_{\mathbf{n}}-\mathbf{n}}t^{\mathbf{n}}=t^{[(s+1)k]_{\mathbf{n}}}b=\sigma^{[(s+1)k]_{\mathbf{n}}}(b)t^{[(s+1)k]_{\mathbf{n}}},$
which gives the desired expression for $L(\mathsf{z},s+1),$ as required.
We now assume that (3.3) holds, and show that this implies power-associativity. By induction it suffices to prove that for all $\ell,s\in\mathbb{N}$ we have
$L(\mathsf{z},s)L(\mathsf{z},\ell)=L(\mathsf{z},s+\ell).$
This holds for $s=1$ and every $\ell\in\mathbb{N}$. We compute
$L$ $(\mathsf{z},s)\;L(\mathsf{z},\ell)=$
$=N_{s}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-i\mathbf{n}}(b)t^{[sk]_{\mathbf{n}}}\cdot N_{\ell}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{\ell k}{\mathbf{n}}\rfloor}\sigma^{\ell k-i\mathbf{n}}(b)t^{[\ell k]_{\mathbf{n}}}$
$=N_{s}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-i\mathbf{n}}(b)\cdot\sigma^{[sk]_{\mathbf{n}}}\left(N_{\ell}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{\ell k}{\mathbf{n}}\rfloor}\sigma^{\ell k-i\mathbf{n}}(b)\right)\cdot t^{[sk]_{\mathbf{n}}+[\ell k]_{\mathbf{n}}}$
By Lemma 3.2, our hypothesis (3.3) implies
$\sigma^{[sk]_{\mathbf{n}}}(N_{\ell}^{\sigma^{k}}(\alpha))\prod_{j=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-j\mathbf{n}}(b)=\sigma^{sk}(N_{\ell}^{\sigma^{k}}(\alpha))\prod_{j=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-j\mathbf{n}+\ell k}(b).$
Thus
$L$ $(\mathsf{z},s)\;L(\mathsf{z},\ell)=$
$=N_{s}^{\sigma^{k}}(\alpha)\left(\prod_{i=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-i\mathbf{n}}(b)\cdot\sigma^{[sk]_{\mathbf{n}}}(N_{\ell}^{\sigma^{k}}(\alpha))\right)\sigma^{[sk]_{\mathbf{n}}}\left(\prod_{i=1}^{\lfloor\frac{\ell k}{\mathbf{n}}\rfloor}\sigma^{\ell k-i\mathbf{n}}(b)\right)\cdot t^{[sk]_{\mathbf{n}}+[\ell k]_{\mathbf{n}}}$
$=N_{s}^{\sigma^{k}}(\alpha)\left(\sigma^{sk}(N_{\ell}^{\sigma^{k}}(\alpha))\prod_{j=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-j\mathbf{n}+\ell k}(b)\right)\prod_{i=1}^{\lfloor\frac{\ell k}{\mathbf{n}}\rfloor}\sigma^{[sk]_{\mathbf{n}}+\ell k-i\mathbf{n}}(b)\cdot t^{[sk]_{\mathbf{n}}+[\ell k]_{\mathbf{n}}}.$
Since $sk]_{\mathbf{n}}=sk-\lfloor\frac{sk}{\mathbf{n}}\rfloor\mathbf{n}$, we may reindex the last product to yield
$=N_{s+\ell}^{\sigma^{k}}(\alpha)\prod_{j=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-j\mathbf{n}+\ell k}(b)\prod_{j=\lfloor\frac{sk}{\mathbf{n}}\rfloor+1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor+\lfloor\frac{\ell k}{\mathbf{n}}\rfloor}\sigma^{sk-j\mathbf{n}+\ell k}(b)\cdot t^{[sk]_{\mathbf{n}}+[\ell k]_{\mathbf{n}}}$
$=N_{s+\ell}^{\sigma^{k}}(\alpha)\prod_{j=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor+\lfloor\frac{\ell k}{\mathbf{n}}\rfloor}\sigma^{(s+\ell)k-j\mathbf{n}}(b)\cdot t^{[sk]_{\mathbf{n}}+[\ell k]_{\mathbf{n}}}.$
If $sk]_{\mathbf{n}}+[\ell k]_{\mathbf{n}}=[(s+\ell)k]_{\mathbf{n}}$ then this is simply
$N_{s+\ell}^{\sigma^{k}}(\alpha)\prod_{j=1}^{\lfloor\frac{(s+\ell)k}{\mathbf{n}}\rfloor}\sigma^{(s+\ell)k-j\mathbf{n}}(b)\cdot t^{[(s+\ell)k]_{\mathbf{n}}}=L(\mathsf{z},s+\ell).$
Otherwise, we must have
$\lfloor\frac{(s+\ell)k}{\mathbf{n}}\rfloor=\lfloor\frac{sk}{\mathbf{n}}\rfloor+\lfloor\frac{\ell k}{\mathbf{n}}\rfloor+1,$
whence $sk]_{\mathbf{n}}+[\ell k]_{\mathbf{n}}=[(s+\ell)k]_{\mathbf{n}}+\mathbf{n}.$ Since $(s+\ell)k]_{\mathbf{n}}=(s+\ell)k-\lfloor\frac{(s+\ell)k}{\mathbf{n}}\rfloor\mathbf{n}$ we have
$t^{[sk]_{\mathbf{n}}+[\ell k]_{\mathbf{n}}}=t^{[(s+\ell)k]_{\mathbf{n}}+\mathbf{n}}=\sigma^{(s+\ell)k-\lfloor\frac{(s+\ell)k}{\mathbf{n}}\rfloor\mathbf{n}}(b)t^{[(s+\ell)k]_{\mathbf{n}}},$
which again yields $L(\mathsf{z},s)L(\mathsf{z},\ell)=L(\mathsf{z},s+\ell).$ Thus the desired equality holds for all $s,\ell\in\mathbb{N}$, and $\mathsf{z}$ is power associative.
We now prove that if $\mathsf{z}\mathsf{z}^{r}=\mathsf{z}^{r}\mathsf{z}$, for $r>0$ the least integer such that $rk\geq\mathbf{n}$, then (3.3) holds. We begin by noting that $\mathsf{z}^{r}$ is well-defined, since for all $s,\ell\geq 1$ such that $s+\ell=r$,
$\mathsf{z}^{s}=N_{s}^{\sigma^{k}}(\alpha)t^{sk},\quad\text{and}\quad\mathsf{z}^{\ell}=N_{\ell}^{\sigma^{k}}(\alpha)t^{\ell k},$
by the associativity of powers less than $\mathbf{n}$ in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$. Thus we readily compute as above that
$\mathsf{z}^{r}=\mathsf{z}^{s}\mathsf{z}^{\ell}=N_{s}^{\sigma^{k}}(\alpha)\sigma^{sk}(N_{\ell}^{\sigma^{k}}(\alpha))t^{(s+\ell)k}=N_{r}^{\sigma^{k}}(\alpha)\sigma^{rk-\mathbf{n}}(b)t^{rk-\mathbf{n}},$
independently of the choice of association, and this is equal to $L(\mathsf{z},r)$.
Now by our choice of $r$ (and $k$), we have $\mathbf{n}<(r+1)k<2\mathbf{n}$ so that $\lfloor(r+1)/\mathbf{n}\rfloor=1$ and $(r+1)k]_{\mathbf{n}}=(r+1)k-\mathbf{n}$. Thus (3.5) simplifies to
$\mathsf{z}\mathsf{z}^{r}=L(\mathsf{z},r+1)=N_{r+1}^{\sigma^{k}}(\alpha)\sigma^{(r+1)k-\mathbf{n}}(b)t^{(r+1)k-\mathbf{n}}$
whereas
$\mathsf{z}^{r}\mathsf{z}=N_{r}^{\sigma^{k}}(\alpha)\sigma^{rk-\mathbf{n}}(b)t^{rk-\mathbf{n}}\alpha t^{k}=N_{r}^{\sigma^{k}}(\alpha)\sigma^{rk-\mathbf{n}}(\alpha)\sigma^{rk-\mathbf{n}}(b)t^{(r+1)k-\mathbf{n}}.$
If these are equal then we may equate these coefficients of $t^{(r+1)k-\mathbf{n}}$. If we divide both expressions by $N_{r}^{\sigma^{k}}(\alpha)$ and simplify, we obtain the equality
$\sigma^{rk}(\alpha)\sigma^{rk-\mathbf{n}}(\sigma^{k}(b))=\sigma^{rk-\mathbf{n}}(\alpha)\sigma^{rk-\mathbf{n}}(b),$
which, upon applying the automorphism $\sigma^{\mathbf{n}-rk}$ to both sides, is simply (3.3). ∎
In particular, Theorem 3.3 shows that the condition (3.3) is both necessary and sufficient for full power associativity of the element $\alpha t^{k}$ in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$.
Corollary 3.4.
Let $k\in\mathbb{N}$ and $\alpha\in S$ not be a zero divisor. Let $\tau$ be an automorphism of $S$ commuting with the order-$\mathbf{m}$ automorphism $\sigma$. Then the map
$G=G_{\tau,\alpha,k}:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
given by $G(a)=\tau(a)$ and $G(t)=\alpha t^{k}$ induces a homomorphism of $S_{0}^{\tau}$-algebras if and only if
$b\in S_{0}^{\times},\quad k\equiv 1\mod\mathbf{m},\quad\text{and}\quad\mathbf{m}|\mathbf{n}.$
Proof.
By Lemma 3.1, the condition $k\equiv 1\mod\mathbf{m}$ ensures that $G(ta)=G(\sigma(a)t)$. By Theorem 3.3, if $G$ is well-defined then $G(t^{s})=\lambda t^{[sk]_{\mathbf{n}}}$ for some $\lambda\in S^{\times}$. Thus the requirement that $G(t^{s})G(a)=G(\sigma^{s}(a))G(t^{s})$ for all $a\in S$ and all $s\in\mathbb{N}$ is equivalent to
$\sigma^{[sk]_{\mathbf{n}}}(\tau(a))=\tau(\sigma^{s}(a)).$
Since $\tau$ commutes with $\sigma$ this holds if and only if $s-[sk]_{\mathbf{n}}\equiv 0\mod\mathbf{m}$; since $k\equiv 1\mod\mathbf{m}$ this implies $\mathbf{m}|\mathbf{n}$. Under this hypothesis, the condition $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)$ simplifies to $\sigma(b)=b$ or $b\in S_{0}^{\times}$, and thus $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is associative. Consequently the function $G=G_{\tau,\alpha,k}$ descends to a well-defined homomorphism $G:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$. ∎
Note that the homomorphism $G$ in Corollary 3.4 is a surjection when additionally $\alpha\in S^{\times}$ and $\gcd(k,\mathbf{n})=1$.
These conditions are quite restrictive, reflecting the constrained nature of these nonassociative algebras. For example, one deduces that even with $k=1$, the map $G_{\tau,\alpha,1}$ is a homomorphism from $S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ only when $b\in S_{0}^{\times}$ and $\mathbf{m}|\mathbf{n}$. This is in fact to be expected: $S[t;\sigma]$ is associative whereas $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is associative if and only if these two strong conditions hold [Pet67, (15)].
4. Monomial homomorphisms between Petit algebras
We now turn to the question of when there exist monomial homomorphisms $G_{\tau,\alpha,k}$ from a Petit algebra $S[t;\sigma]/S[t;\sigma]f(t)$, where $f$ has degree $\mathbf{n}$, into the algebra $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$.
It is not evident that $G$ is forced to be the reduction of a homomorphism $\tilde{G}:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ as in Corollary 3.4. Namely, in $S[t;\sigma]/S[t;\sigma]f(t)$, the associativity of all powers of $t$ is a strong condition [Pet67, p. 13-06]: it is equivalent to $t^{\mathbf{n}}t=tt^{\mathbf{n}}$ and to $f(t)t\in S[t;\sigma]f(t)$. In particular, if all powers of $t$ are associative in $S[t;\sigma]/S[t;\sigma]f(t)$, then they are contained in the right nucleus of $S[t;\sigma]/S[t;\sigma]f(t)$, which is an associative subalgebra, and thus it is only in this case that Theorem 3.3 applies directly, as in the following example.
Example 4.1.
Suppose $f(t)\in S_{0}[t]$. Then the associative algebra $S_{0}[t]/\langle f(t)\rangle$ is a subalgebra of $S[t;\sigma]/S[t;\sigma]f(t)$ containing all powers of $t$. It thus follows from Theorem 3.3 that if $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a monomial homomorphism, then we must have
$\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)$
and by (3.4), for every $s\in\mathbb{N}$, the expression $G(t^{s})=G(t)^{s}$ is equal to
(4.1) $N_{s}^{\sigma^{k}}(\alpha)\prod_{i=1}^{\lfloor\frac{sk}{\mathbf{n}}\rfloor}\sigma^{sk-i\mathbf{n}}(b)t^{[sk]_{\mathbf{n}}}.$
When instead $S[t;\sigma]/S[t;\sigma]f(t)$ is proper nonassociative, then $t^{s}$ is only well-defined for $0\leq s<\mathbf{n}$, and the full power associativity of Theorem 3.3 is not required.
To illustrate this, we derive the following generalization of Theorem 2.2, about the existence of monomial homomorphisms of degree $1$, between Petit algebras of our chosen form.
Proposition 4.2.
Let $\tau\in\operatorname{Aut}(S)$ commute with $\sigma$ and suppose $\alpha\in S$ is not a zero divisor. Then the map
$G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
induces an homomorphism if and only if
$\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b,$
and this is an isomorphism if and only if $\alpha\in S^{\times}$.
In particular, there is no requirement on the relationship between $\mathbf{n}$ and $\mathbf{m}$.
Proof.
Write $G=G_{\tau,\alpha}$; then $G(t)=\alpha t$ and so for all $1\leq s<\mathbf{n}$ we unambiguously have $G(t^{s})=N_{s}^{\sigma}(\alpha)t^{s}$ since these powers are associative in both rings. Thus since $G$ acts by $\tau$ on $S$, $G$ extends to an $S_{0}^{\tau}$-linear isomorphism of the underlying $S_{0}^{\tau}$-modules.
Furthermore, for all $c\in S$ and $0<s<\mathbf{n}$ we have $t^{s}c=\sigma^{s}(c)t^{s}$. We find
$G(t^{s})G(c)=N_{s}^{\sigma}(\alpha)t^{s}\tau(c)=N_{s}^{\sigma}(\alpha)\sigma^{s}(\tau(c))t^{s}=\tau(\sigma^{s}(c))N_{s}^{\sigma}(\alpha)t^{s}=G(\sigma^{s}(c))G(t^{s})$
and thus this identity always holds.
We now verify that $G(pq)=G(p)G(q)$ for all $p,q\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$. By the above, it suffices to choose arbitrary $1\leq s,\ell<\mathbf{n}$ and prove this for $p=t^{s}$, $q=t^{\ell}$.
Consider the product $t^{s}t^{\ell}$ in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$. If $s+\ell<\mathbf{n}$ the result is $t^{s+\ell}$ and we correspondingly have
$G(t^{s})G(t^{\ell})=N_{s}^{\sigma}(\alpha)t^{s}N_{\ell}^{\sigma}(\alpha)t^{\ell}=N_{s}^{\sigma}(\alpha)\sigma^{s}(N_{\ell}^{\sigma}(\alpha))t^{s+\ell}=N_{s+\ell}^{\sigma}(\alpha)t^{s+\ell}$
which is equal to $G(t^{s+\ell})$. If $s+\ell\geq\mathbf{n}$, however, then we have $t^{s}t^{\ell}=t^{s+\ell-\mathbf{n}}a=\sigma^{s+\ell-\mathbf{n}}(a)t^{s+\ell-\mathbf{n}}$ whose image under $G$ is
$G(t^{s}t^{\ell})=\tau(\sigma^{s+\ell-\mathbf{n}}(a))N_{s+\ell-\mathbf{n}}^{\sigma}(\alpha)t^{s+\ell-\mathbf{n}}$
while
$G(t^{s})G(t^{\ell})=N_{s+\ell}^{\sigma}(\alpha)t^{s+\ell-\mathbf{n}}b=N_{s+\ell}^{\sigma}(\alpha)\sigma^{s+\ell-\mathbf{n}}(b)t^{s+\ell-\mathbf{n}}.$
Since $N_{s+\ell}^{\sigma}(\alpha)=N_{s+\ell-\mathbf{n}}^{\sigma}(\alpha)\sigma^{s+\ell-\mathbf{n}}(N_{\mathbf{n}}^{\sigma}(\alpha))$ by (2.1) and $N^{\sigma}_{s+\ell-\mathbf{n}}(\alpha)$ is assumed to be cancellable, we deduce that these two expressions are equal whenever
$\tau(\sigma^{s+\ell-\mathbf{n}}(a))=\sigma^{s+\ell-\mathbf{n}}(N_{\mathbf{n}}^{\sigma}(\alpha))\sigma^{s+\ell-\mathbf{n}}(b).$
Therefore, since $\tau$ and $\sigma$ commute, our map is a homomorphism of $S_{0}^{\tau}$-algebras if and only if $\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b$. It is invertible exactly when $\alpha^{-1}\in S$. ∎
For example, when $\mathbf{m}=\mathbf{n}$, $S=K$ is a Galois field extension of $K_{0}$ and $\tau\in{\rm Gal}(K/K_{0})$, we recover the condition that $\tau(a)\in bN_{K/K_{0}}(K^{\times})$ from [BP18, Corollary 32].
We now turn to the case that the degree $k$ of the monomial homomorphism $G_{\tau,\alpha,k}$ is strictly greater than $1$. We begin with a lemma that is a version of Lemma 3.1 for the present setting.
Lemma 4.3.
Suppose $f,g\in S[t;\sigma]$ are monic polynomials of degree $\mathbf{n}$. If for some cancellable $\alpha\in S$ and degree $k>1$, the map $G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma]g(t)$ is a monomial homomorphism, then $k\equiv 1\mod\mathbf{m}$; if it is an isomorphism then $\gcd(k,\mathbf{n})=1$ and $\alpha\in S^{\times}$.
Proof.
If $G=G_{\tau,\alpha,k}$ is a homomorphism, then the identity $G(td)=G(\sigma(d)t)$ implies $\alpha t^{k}\tau(d)=\tau(\sigma(d))\alpha t^{k}$ for all $d\in S$, whence we must have
$\sigma^{k}(\tau(d))=\tau(\sigma(d))$
for all $d\in S$. Thus since $\tau$ and $\sigma$ commute, $G(td)=G(\sigma(d)t)$ is equivalent to $k\equiv 1\mod\mathbf{m}$.
If $G$ is to be an isomorphism, there must exist an inverse homomorphism $H$. Evidently $H(d)=\tau^{-1}(d)$ for all $d\in S$. Since $G$ is monomial, so is its inverse, and thus we may assume $H(t)=\beta t^{\ell}$ for some $\beta\in S^{\times}$ and $1\leq\ell<\mathbf{n}$. Since $H(G(t))=t$ we must have $\ell k\equiv 1\mod\mathbf{n}$ and in particular $\gcd(k,\mathbf{n})=1$. Moreover, if $\alpha$ were not invertible, then $t^{k}$ could not lie in the image of $G$. ∎
Theorem 4.4.
Suppose $\tau\in\operatorname{Aut}(S)$ commutes with $\sigma$ and $\alpha\in S$ is cancellable. Then for any $2\leq k<\mathbf{n}$, the map
$G=G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
is a homomorphism of rings if and only if all of the following conditions hold:
  1. (1)
    $k\equiv 1\mod\mathbf{m}$;
  2. (2)
    $\mathbf{m}|\mathbf{n}$;
  3. (3)
    $a,b\in S_{0}$;
  4. (4)
    $(N_{S/S_{0}}(\alpha))^{\mathbf{n}/\mathbf{m}}b^{k}=\tau(a)$.
It is an isomorphism if and only if, in addition, we have
  1. (5)
    $\gcd(k,\mathbf{n})=1$ and $\alpha\in S^{\times}$.
Proof.
Let $\tau\in\operatorname{Aut}(S)$ commute with $\sigma$. Fix $\alpha\in S^{\times}$ and $2\leq k<\mathbf{n}$. Suppose first that the map $G=G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ defined by the rule
(4.2) $G\left(\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\right)=\sum_{i=0}^{\mathbf{n}-1}\tau(a_{i})(\alpha t^{k})^{i}$
is a well-defined homomorphism.
Conditions (1) and (5) follow from Lemma 4.3.
For $G$ to be well-defined, we require for all $0<s<\mathbf{n}$ that $G(t^{s})=G(t)^{s}$ is well-defined. Since $k>1$, the minimal value of $r$ such that $rk\geq\mathbf{n}$ satisfies $0<r<\mathbf{n}$ and thus by Theorem 3.3, we conclude that $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)$. Conversely, Theorem 3.3 further assures us that this hypothesis implies all powers of $\alpha t^{k}$ are well-defined so (4.2) gives a well-defined map.
Next, we note from Theorem 3.3 that for any $0<s<\mathbf{n}$, the element $G(t^{s})=G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ will be a monomial of degree $sk]_{\mathbf{n}}$, the residue of $sk$ mod $\mathbf{n}$. In order for $G$ to further preserve the relation $t^{s}d=\sigma^{s}(d)t^{s}$ for all $d\in S$ we require $G(t^{s})\tau(d)=\tau(\sigma^{s}(d))G(t^{s})$, which will hold if and only if
$\sigma^{[sk]_{\mathbf{n}}}(\tau(d))=\tau(\sigma^{s}(d))$
for all $d\in S$. Since $k\equiv 1\mod\mathbf{m}$, and $\tau\sigma=\sigma\tau$, this is equivalent to the requirement that $\sigma^{sk}=\sigma^{[sk]_{\mathbf{n}}}$ for all $0<s<\mathbf{n}$, which holds if and only if $\sigma^{\mathbf{n}}=1$. Thus a second necessary condition is (2), that $\mathbf{m}|\mathbf{n}$, and it in return implies $G(t^{s})\tau(d)=\tau(\sigma^{s}(d))G(t^{s})$ for all $0<s<\mathbf{n}$ and $d\in S$.
Since $\mathbf{m}|\mathbf{n}$ and $k\equiv 1\mod\mathbf{m}$, the relation $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)$ simplifies to $\sigma(b)=b$, or $b\in S_{0}$, yielding part of the condition (3).
Next, $G$ must satisfy $G(t^{\mathbf{n}}-a)=0$. By the above, we may evaluate $G(t^{\mathbf{n}})=G(t)^{\mathbf{n}}$ using the formula (3.4). With $\sigma^{k}=\sigma$ and $\sigma^{\mathbf{n}}=1$, this yields
$G(t)^{\mathbf{n}}=N_{\mathbf{n}}^{\sigma}(\alpha)\left(\prod_{i=1}^{k}b\right)t^{0}=(N^{\sigma}_{\mathbf{m}}(\alpha))^{\mathbf{n}/\mathbf{m}}b^{k}=(N_{S/S_{0}}(\alpha))^{\mathbf{n}/\mathbf{m}}b^{k}.$
Now since $G(t^{\mathbf{n}})=G(a)$, this is equivalent to the condition (4)
$(N_{S/S_{0}}(\alpha))^{\mathbf{n}/\mathbf{m}}b^{k}=\tau(a).$
In particular $\tau(a)\in S_{0}$, so $a\in S_{0}$, which is the rest of condition (3).
Note that if $G$ is an isomorphism, its inverse map must be a homomorphism. If $G^{-1}(t)=\beta t^{\ell}$, then $\ell\equiv 1\mod\mathbf{m}$; in fact, $\ell$ is simply the inverse of $k$ mod $\mathbf{n}$ (as guaranteed by condition (5)).
Finally, if $k,\mathbf{n},\mathbf{m},\alpha,a,b,\tau$ satisfy the first four conditions, then the $S_{0}^{\tau}$-linear map $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ given by (4.2) is a well-defined homomorphism of $S_{0}^{\tau}$-modules, and respects the relations $G(t^{\mathbf{n}})=G(a)$ and $G(t^{s}u)=G(\sigma^{s}(u)t^{s})$ for all $0<s<\mathbf{n}$ and all $u\in S$. Therefore in particular it satisfies $G(p)G(q)=G(pq)$ and $G(p+q)=G(p)+G(q)$ for all $p,q\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$, which implies it is an homomorphism of $S_{0}^{\tau}$-algebras. The final condition ensures it is an isomorphism. ∎
An immediate consequence is the realization that there are very few monomial homomorphisms of degree greater than one in this setting.
Corollary 4.5.
Suppose $\mathbf{m}\nmid\mathbf{n}$, or that one of $a$ or $b$ is not in $S_{0}$. Then any Hamming-weight-preserving homomorphism
$G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
will be of degree one, that is, will be of the form $G=G_{\tau,\alpha}$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$, and some $\alpha\in S$ that is not a zero divisor.
Proof.
If $G$ is Hamming weight preserving, then it must send $t$ to an element of Hamming weight $1$, hence it is monomial. Further, the coefficient $\alpha$ in this case cannot be a zero divisor, or else there would be some $c\in S$ for which $G(ct)=0$. By Theorem 4.4, the degree of $G$ cannot be greater than one. ∎
When $a,b\in S_{0}$, on the other hand, Theorem 4.4 gives a plethora of monomial homomorphisms $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ of degree $k>1$.
Example 4.6.
Suppose $\mathbf{m}=2,\mathbf{n}=4$ and choose $k=\ell=3$. For any $S$, and any $r=\alpha\sigma(\alpha)\in N_{S/S_{0}}(S^{\times})$, the map $G=G_{\operatorname{id},\alpha,3}$ induces an isomorphism
$S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-r^{2}b^{3})\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
for any $b\in S^{\times}$.
5. Applications
In Section 4, we classified all monomial homomorphisms between Petit algebras of the form $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$. Since the left ideals of these algebras give rise to skew constacyclic codes, we have classified the Hamming-weight preserving maps, or isometries, between their corresponding skew constacyclic codes.
5.1. Isometric constacyclic codes
In [OaNA25], the authors explore Hamming-weight preserving isomorphisms between nonassociative algebras of the form
$K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$
over a finite field $K=\mathbb{F}_{p^{r}}$ with respect to a Frobenius automorphism $\sigma(a)=a^{p^{s}}$ for some $s|r$, as an important step towards classifying skew constacyclic codes of length $\mathbf{n}$ with identical performance parameters. This is necessary to avoid duplicating already existing codes. To this end, they proposed notions of $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry, respectively.
We say two rings $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are $(\mathbf{n},\sigma)$-equivalent if they are isomorphic via an isomorphism of the form $G_{\operatorname{id},\alpha}$ for some $\alpha\in K^{\times}$, and $(\mathbf{n},\sigma)$-isometric if they are isomorphic via an isomorphism of the form $G_{\operatorname{id},\alpha,k}$ for some $\alpha\in K^{\times}$ and $k\geq 1$.
By Corollary 4.5, we conclude that these notions coincide whenever $a$ or $b$ lies in $K_{0}=\mathbb{F}_{p^{s}}$, and whenever the order of $\sigma$ (which is $r/s$) does not divide $\mathbf{n}$.
Inspired by their work, we now propose a more complete notion of equivalence for such rings.
Definition 5.1.
Two rings $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ and $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ are called isometric if there exists a monomial isomorphism $G_{\tau,\alpha,k}$ between them for some $\alpha\in S^{\times}$ and $1\leq k<\mathbf{n}$. As a special case, we say they are equivalent if they are isomorphic via an isomorphism of the form $G_{\tau,\alpha}$ for some $\alpha\in S^{\times}$ and some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$.
Thus if two rings are $(\mathbf{n},\sigma)$-isometric, or $(\mathbf{n},\sigma)$-equivalent, then they are isometric, and if they are $(\mathbf{n},\sigma)$-equivalent then they are equivalent. Moreover, this definition captures all of the possible Hamming-weight preserving isomorphisms between these algebras, making it the correct notion of equivalence between codes.
5.2. A surprising consequence
The surprising scarcity of monomial isomorphisms (as shown here in Theorem 4.4) sheds light on an error in [OaNA25, Theorem 4], where four statements are said to be equivalent. In fact, trying to uncover the truth of this theorem was the impetus for our study.
Set $K=\mathbb{F}_{q}=\mathbb{F}_{p^{r}}$ with primitive element $\xi$, and let $\sigma(x)=x^{p^{s}}$. (Note that the order of $\sigma$ is $r/s$, not $s$, as stated.) Then $N_{\mathbf{n}}^{\sigma}(x)=x^{[\mathbf{n}]_{s}}$ for all $x\in K$ where
$\mathbf{n}]_{s}=\frac{p^{s\mathbf{n}}-1}{p^{s}-1}=p^{s(\mathbf{n}-1)}+p^{s(\mathbf{n}-2)}+\cdots+p^{s}+1.$
Then the subgroup of $K^{\times}$ consisting of all possible values of $N_{\mathbf{n}}^{\sigma}(\alpha)$ is the one generated by $N_{\mathbf{n}}^{\sigma}(\xi)=\xi^{[\mathbf{n}]_{s}}$.
In [OaNA25], $a,b\in S^{\times}$ are called $(\mathbf{n},\sigma)$-equivalent if $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are $(\mathbf{n},\sigma)$-equivalent and $(\mathbf{n},\sigma)$-isometric if $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are $(\mathbf{n},\sigma)$-isometric. Thus briefly, [OaNA25, Theorem 4] asserted the equivalence of:
(1):
$K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are $(\mathbf{n},\sigma)$-isometric;
(2) and (3):
$\langle a,N_{\mathbf{n}}^{\sigma}(\xi)\rangle=\langle b,N_{\mathbf{n}}^{\sigma}(\xi)\rangle$;
(4):
$K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b^{k})$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ are $(\mathbf{n},\sigma)$-equivalent, where $k$ is the degree of the isometry in (1). [^1] The actual conditions on $k$ specified were: $0<k<[\mathbf{m}]_{s}$, $\gcd(k,[\mathbf{m}]_{s})$.
Note that (1) implies (2) and (3) by Proposition 4.2 (for monomial homomorphisms of degree $1$) and Theorem 4.4 (for monomial homomorphisms of degree $k>1$, when they exist). However, the converse implication fails, as illustrated in the following counterexample.
Example 5.2.
Choose the parameters $p,r,s,\mathbf{n}$ to satisfy that $(p^{r}-1)|[\mathbf{n}]_{s}$. For example, one may take $p=3$, $r=2$, $s=1$ and $\mathbf{n}=4$, so that $\mathbf{m}=2$ and $p^{r}-1=8$ (meaning $\xi^{8}=1$) and $\mathbf{n}]_{s}=40$. Then $\langle\xi^{[\mathbf{n}]_{s}}\rangle=\{1\}$ and in this case there are very few isomorphisms between these rings.
We now choose $a$ so that $a$ and $b=a^{k}$ are two primitive elements of the field that are not Galois conjugates of one another; together with the triviality $N_{\mathbf{n}}^{\sigma}(\alpha)$ for all $\alpha$, this guarantees that $\tau(a)\notin bN_{\mathbf{n}}^{\sigma}(\alpha)$.
For example, let $a=\xi$ and $b=a^{k}$ for some $2\leq k<p^{r}-1$ such that $k\neq p^{\ell}$ for any $\ell$ (ensuring $b\neq\tau(a)$ for any $\tau\in\operatorname{Gal}(K/\mathbb{F}_{p})$) and $\gcd(k,p^{2}-1)=1$ (ensuring $b$ is another primitive element). Then we have
$\langle a,\xi^{[\mathbf{n}]_{s}}\rangle=K^{\times}=\langle b,\xi^{[\mathbf{n}]_{s}}\rangle$
so that (2) holds. Now we claim (1) fails.
Since $N_{\mathbf{n}}^{\sigma}(\alpha)=1$ for all $\alpha$, Proposition 4.2 implies that if there exists an monomial isomorphism of degree one $G_{\tau,\alpha}:K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)\to K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ we would have $\tau(a)=b$, which is false. Moreover, since $a\notin K_{0}$, condition Theorem 4.4[(3)] fails, so there are no monomial homomorphisms of degree $k$ (for any $k>1$). Thus $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are not isometric, and (1) fails.
The issue is that the condition for isomorphism, that is, that $a\in b\langle\xi^{[\mathbf{n}]_{s}}\rangle$, is strictly weaker than (2), even if one would consider the full class of isomorphisms of the form $G_{\tau,\alpha,k}$.
Further: taking $k=1$ we see that (1) implies (4), but again our results show that the converse is false. We offer the same counterexample.
Example 5.3.
We return to the setting of Example 5.2, where $b=a^{k}$. Then $\operatorname{id}:K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a^{k})\to K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism satisfying (4), but we have already shown that $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are not isometric, so again (1) fails.
These errors stemmed from an expectation of the existence of many monomial homomorphisms of degree other than one. They invalidate the proofs of [OaNA25, Theorem 5, Remark 3, Corollary 1, 3, Remark 4] and the proposed classification of skew constacyclic codes up to $(\mathbf{n},\sigma)$-isometry in [OaNA25, Sections 5].
On the bright side, our results show that in many cases, the notions of $(\mathbf{n},\sigma)$-isometry and $(\mathbf{n},\sigma)$-equivalence coincide (as do our more general notions of isometry and equivalence), so that employing $(\mathbf{n},\sigma)$-isometry as an equivalence relation to partition skew constacyclic codes will in many cases suprisingly not lead to a finer partition of skew constacyclic codes with the same performance parameters, as we might have all expected initially. This gives considerably more weight to the results obtained in [OaNA25, Sections 6, 7].
5.3. An interesting correction
Let us now formulate which results do hold and generalize them immediately to the setting of unital commutative rings $S$ and allow $\tau$ to be a nontrivial automorphism.
Lemma 5.4.
Let $\sigma\in{\rm Aut}(S)$ have order $\mathbf{m}$, fixed field $S_{0}$, and let $a,b,\alpha\in S^{\times}$. If there is a Hamming-weight preserving isomorphism $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ with $G|_{S}=\tau$, where $\tau\in\operatorname{Aut}(S)$ commutes with $\sigma$, then
$\langle b,N_{\mathbf{n}}^{\sigma}(S^{\times})\rangle=\langle\tau(a),N_{\mathbf{n}}^{\sigma}(S^{\times})\rangle.$
In particular, the subgroups in $S^{\times}/N_{\mathbf{n}}^{\sigma}(S^{\times})$ generated by $\tau(a)$ and $b$ are the same.
Proof.
Suppose there exists a Hamming weight preserving algebra isomorphism given by $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ such that $G|_{S}=\tau$ and $G(t)=\alpha t^{k}$ for some $\alpha\in S^{\times}$, $0<k\leq\mathbf{n}-1$.
We now distinguish two cases, employing Theorem 4.4:
If $a\in S\setminus S_{0}$ or $a\in S_{0}^{\times}$ but $\mathbf{m}\nmid\mathbf{n}$ then $k=1$ and so $\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b$ by Proposition 4.2, hence $\tau(a)$ lies in the subgroup of $S^{\times}$ that is generated by $b$ and $N_{\mathbf{n}}^{\sigma}(S^{\times})$. This means $\langle\tau(a),N_{\mathbf{n}}^{\sigma}(S^{\times})\rangle\subset\langle b,N_{\mathbf{n}}^{\sigma}(S^{\times})\rangle$.
If $a,b\in S_{0}^{\times}$ and $\mathbf{m}|\mathbf{n}$ then $k>1$ is possible, and we have by Theorem 4.4 that $k\equiv 1\mod\mathbf{m}$ and $\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b^{k}$. This implies that $gcd(k,\mathbf{n})=1$ and thus $\tau(a)$ lies in the subgroup of $S^{\times}$ that is generated by $b$ and $N_{\mathbf{n}}^{\sigma}(S^{\times})$. This means $\langle\tau(a),N_{\mathbf{n}}^{\sigma}(S^{\times})\rangle\subset\langle b,N_{\mathbf{n}}^{\sigma}(S^{\times})\rangle$.
In both cases, the inverse isomorphism is given by $G_{\tau^{-1},\beta,\ell}$ for some $\beta\in S^{\times}$ and $0<\ell\leq\mathbf{n}-1$, so applying the same arguments yields $\langle\tau^{-1}(b),N_{\mathbf{n}}^{\sigma}(S^{\times})\rangle\subset\langle a,N_{\mathbf{n}}^{\sigma}(S^{\times})\rangle.$ Since $\tau$ commutes with $\sigma$, $\tau(N_{\mathbf{n}}^{\sigma}(S^{\times}))=N_{\mathbf{n}}^{\sigma}(S^{\times})$ and thus we obtain the desired equality. ∎
As pointed out in Example 5.2, the statement in the lemma is not an equivalence, in general, not even over finite fields.
Proposition 5.5.
Let $\sigma\in{\rm Aut}(S)$ have order $\mathbf{m}$, fixed field $S_{0}$, and let $a,b,\alpha\in S^{\times}$. Let $\tau\in\operatorname{Aut}(S)$ be any automorphism that commutes with $\sigma$.
  1. (a)
    Suppose $G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism with $k>1$ for some $\alpha\in S^{\times}$. Then $k\equiv 1\mod\mathbf{m}$, $\gcd(k,\mathbf{n})=1$, and $G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b^{k})$ is also an isomorphism.
  2. (b)
    Then $G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism if and only if $G_{id,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-\tau(a))\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism.
  3. (c)
    Also, for any $a\in\mathrm{Fix}(\tau)=\{b\in S:\tau(b)=b\}$, we have $G_{\operatorname{id},\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism if and only if $G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism.
Proof.
(a) Suppose there exists an isomorphism
$G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
for some $\alpha\in S^{\times}$, $1<k<\mathbf{n}$. By Theorem 4.4, since $k>1$, we must have $a,b\in S_{0}^{\times}$ and $\mathbf{m}|\mathbf{n}$. Moreover, by Theorem 4.4 we know $k\equiv 1\mod\mathbf{m}$, $\gcd(k,\mathbf{n})=1$, and $\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b^{k}$. Therefore by Proposition 4.2,
$G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b^{k})$
is an isomorphism.
(b) By Proposition 4.2, $G_{\tau,\alpha}$ is an isomorphism if and only if $\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b$. This is equivalent to
$G_{id,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-\tau(a))\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
being a $S_{0}$-algebra isomorphism.
(c) is a special case of (b). ∎
Let us point out that Proposition 5.5 (b) highlights the advantage of choosing equivalence over $(\mathbf{n},\sigma)$-equivalence: when $\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b$ then the class of skew $(\sigma,a)$-constacyclic codes is equivalent to the class of skew $(\sigma,b)$-constacyclic codes, but so will be the class of $(\sigma,\tau(a))$-constacyclic codes for any $\tau$ commuting with $\sigma$.
It also shows that the class of skew $(\sigma,a)$-constacyclic codes is equivalent to the class of skew $(\sigma,b)$-constacyclic codes via $G_{\tau,\alpha}$, if and only if the class of skew $(\sigma,\tau(a))$-constacyclic codes is $(\mathbf{n},\sigma)$-equivalent to the class of $(b,\sigma)$-constacyclic codes.
Moreover, (a) implies that two classes of skew constacyclic codes that are $(\mathbf{n},\sigma)$-isometric and related to principal left ideals in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$, respectively to principal left ideals in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, will generate two classes of codes in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a^{k})$, respectively, in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, that are $(m,\sigma)$-equivalent.
In [OaNA25, Theorem 4] the authors claim that the number of $(\mathbf{n},\sigma)$-isometry classes is equal to the number of divisors of $\gcd([\mathbf{n}]_{s},p^{r}-1)$, but as this was based on the false equivalence (1) $\iff$ (2), the answer was incorrect.
We derive the answer for the case that $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ is nonassociative, that is, when $(\mathbf{n},\sigma)$-equivalence and $(\mathbf{n},\sigma)$-isometry coincide. Note that we do not limit ourselves to those $a$ such that $t^{n}-a$ is reducible.
Theorem 5.6.
Suppose $K=\mathbb{F}_{p^{r}}$ and $\sigma(x)=x^{p^{s}}$ with $s|r$ so that $\mathbf{m}=r/s$ and $K_{0}=\mathbb{F}_{p^{s}}$. When $a\notin K_{0}$, $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ is nonassociative, and the number of different $(\mathbf{n},\sigma)$-isometry classes is $N$ where
$N=\begin{cases}\gcd([\mathbf{n}]_{s},p^{r}-1)&\text{if $\mathbf{m}\nmid\mathbf{n}$; and}\\ \gcd([\mathbf{n}]_{s},p^{r}-1)-1&\text{if $\mathbf{m}|\mathbf{n}$.}\end{cases}$
Moreover, if $w=\gcd([\mathbf{n}]_{s},p^{r}-1)$ and $t=\gcd([\mathbf{n}]_{s},\frac{p^{r}-1}{p^{s}-1})$ then the number of $b\in aN_{\mathbf{n}}^{\sigma}(\mathbb{F}_{p^{r}}^{\times})$ such that $b\notin\mathbb{F}_{p^{s}}$ is
  • $(p^{r}-1)/w$ if $\mathbf{m}|\mathbf{n}$ and $a\notin\mathbb{F}_{p^{s}}$
  • 0 if $\mathbf{m}|\mathbf{n}$ and $a\in\mathbb{F}_{p^{s}}$
  • $(p^{r}-1)/w$ if $\mathbf{m}\nmid\mathbf{n}$ and $a=\xi^{A}$ such that $t\nmid A$
  • $(p^{r}-1)/w-t$ if $\mathbf{m}\nmid\mathbf{n}$ and $a=\xi^{A}$ and $t|A$.
Note that when $\mathbf{m}|\mathbf{n}$, we have $N_{\mathbf{n}}^{\sigma}(K)\subset K_{0}$, and so if $a\in bN_{\mathbf{n}}^{\sigma}(K)$ and $a\notin K_{0}$ then $b\notin K_{0}$. Thus the second case corresponds to the isometry classes of associative Petit algebras.
Proof.
The $(\mathbf{n},\sigma)$-isometry class of $(\sigma,a)$-constacyclic codes is represented by all $b\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p^{s}}$ such that
$b\in aN_{\mathbf{n}}^{\sigma}(\alpha)$
for some $\alpha\in K^{\times}$, that is, if and only if $a$ and $b$ lie in the same coset of the subgroup $N_{\mathbf{n}}^{\sigma}(K^{\times})$. Let $\xi$ be a primitive element of $K$; it has order $p^{r}-1$. Then $N_{\mathbf{n}}^{\sigma}(K^{\times})$ is generated by $N_{\mathbf{n}}^{\sigma}(\xi)=\xi^{[\mathbf{n}]_{s}}=\xi^{w}$ with $w=\gcd([\mathbf{n}]_{s},p^{r}-1)$, and thus the order of this group is
$\frac{p^{r}-1}{w}.$
The number of equivalence classes of codes is the number of cosets of this group, yielding
$\gcd([\mathbf{n}]_{s},p^{r}-1).$
However, not all of these will be nonassociative so this is an upper bound.
For a more precise count, choose $a\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p^{s}}$ and write $a=\xi^{A}$ for some integer $A$. Then we have $b\in aN_{\mathbf{n}}^{\sigma}(\mathbb{F}_{p^{r}}^{\times})$ if and only if
$b\in\{\xi^{A+kw}\mid 0\leq k<\frac{p^{r}-1}{w}\}.$
Recall that $\xi^{h}\in\mathbb{F}_{q^{s}}$ if and only if $(p^{r}-1)|h(p^{s}-1)$ or $h\equiv 0\mod\frac{p^{r}-1}{p^{s}-1}$. Thus we need to subtract from this coset all choices of $k$ such that
(5.1) $A+kw\equiv 0\mod\frac{p^{r}-1}{p^{s}-1}.$
Recall $\mathbf{m}=r/s$. If $\mathbf{m}|\mathbf{n}$ or $r|s\mathbf{n}$ then $p^{r}-1$ divides $p^{sn}-1$, so that $(p^{r}-1)/(p^{s}-1)$ divides $\gcd([\mathbf{n}]_{s},p^{r}-1)=w$. Thus (5.1) holds if and only if $A\equiv 0\mod(p^{r}-1)/(p^{s}-1)$, that is, if and only if $a\in\mathbb{F}_{p^{s}}$ already. This gives the first two statements.
Now set $t=\gcd(w,\frac{p^{r}-1}{p^{s}-1})=\gcd([\mathbf{n}]_{s},\frac{p^{r}-1}{p^{s}-1})=\gcd(\frac{p^{s\mathbf{n}}-1}{p^{s}-1},\frac{p^{r}-1}{p^{s}-1}).$ So $t$ divides $w$ and $\frac{p^{r}-1}{p^{s}-1}$ so if $t\nmid A$ then (5.1) has no solution, meaning all elements of the coset are outside of $\mathbb{F}_{p^{s}}$.
On the other hand, if $t|A$ write $A=\ell t$; then (5.1) holds for all $k$ such that $\ell+k(w/t)\equiv 0\mod\frac{p^{r}-1}{t(p^{s}-1)}$. Since $\gcd(w/t,\frac{p^{r}-1}{t(p^{s}-1)})=1,$ this has a unique solution modulo $\frac{p^{r}-1}{t(p^{s}-1)}$, hence $t$ solutions modulo $\frac{p^{r}-1}{(p^{s}-1)}$. Therefore $t$ elements of the coset lie in $\mathbb{F}_{p^{s}}$, yielding the final statement.
Therefore, if $\mathbf{m}|\mathbf{n}$ we have exactly $\gcd([\mathbf{n}]_{s},p^{r}-1)-1$ of these $(\mathbf{n},\sigma)$ isometry (or $\mathbf{n},\sigma)$-equivalence) classes, since every coset except one is fully contained in $\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p^{s}}$.
On the other hand, if $\mathbf{m}\nmid\mathbf{n}$ then every coset meets $\mathbb{F}_{p^{s}}\setminus\mathbb{F}_{p^{r}}$, so we have $w$ of these $(\mathbf{n},\sigma)$ isometry (or $\mathbf{n},\sigma)$-equivalence) classes. ∎
Example 5.7.
At one extreme, suppose that $N_{\mathbf{n}}^{\sigma}(S^{\times})=S^{\times}$. Then all $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ are $(\mathbf{n},\sigma)$-equivalent, and in particular all are $(\mathbf{n},\sigma)$-equivalent to $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-1)$, which corresponds to the class of skew cyclic codes. Moreover, equivalence and $(\mathbf{n},\sigma)$-equivalence coincide in this case. When $\mathbf{m}\nmid\mathbf{n}$, then the notions of equivalence, isometry and $(\mathbf{n},\sigma)$-isometry coincide as well. Hence for all $a\in S^{\times}$, the skew $(\sigma,a)$-constacyclic codes of length $\mathbf{n}$ over $S$ are equivalent to the skew cyclic codes of length $\mathbf{n}$.
As a special case, let $K=\mathbb{F}_{p^{r}}$, $\sigma(x)=x^{p^{s}}$ with $s|r$ and $K_{0}=\mathbb{F}_{p^{s}}$. If $\gcd([\mathbf{n}]_{s},p^{r}-1)=1$ then $N_{\mathbf{n}}^{\sigma}(K^{\times})=K^{\times}$ so all $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ are both $(\mathbf{n},\sigma)$-equivalent, as was already observed in [OaNA25, Corollary 2] and [BBB21, Proposition 1].
Example 5.8.
At the other extreme, when $\gcd([\mathbf{n}]_{s},p^{r}-1)=p^{r}-1$ (as in Example 5.2), then $N_{\mathbf{n}}^{\sigma}(K^{\times})=\{1\}$ and so no two distinct $a,b\in K^{\times}$ will be $(\mathbf{n},\sigma)$-equivalent. Thus, there are a plethora of distinct skew constacyclic codes. That said, there always at least $\mathbf{m}$ choices for $\tau$ (corresponding to $\sigma^{\ell}$ for each $1\leq\ell<\mathbf{m}$).
In this example, we deduce that the set $\{a^{p^{v}}\mid 0\leq v<r\}\subset\mathbb{F}_{q^{r}}$ is an equivalence class with $r$ elements. That is, all of the corresponding $(\sigma,a^{p^{v}})$-constacyclic codes are equivalent. Thus there are in fact fewer isometry classes overall. This clearly demonstrates that using our refined notion of equivalence and isometry gives a tighter classification of these classes of codes.
On the other hand, when $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ is associative we can prove a powerful equivalence (inspired by the flavour of [OaNA25, Theorem 4]) that holds for any commutative ring $S$ and any $\tau$ commuting with $\sigma$.
Theorem 5.9.
Suppose that $a,b\in S_{0}^{\times}$ and $\mathbf{m}|\mathbf{n}$. Assume that $\tau$ commutes with $\sigma$. Then the following statements are equivalent for any integer $1\leq k<\mathbf{n}$:
  1. (i)
    $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are isometric via $G_{\tau,\alpha,k}$;
  2. (ii)
    $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b^{k})$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ are equivalent via $G_{\tau,\alpha}$, where $k$ satisfies $k\equiv 1\mod\mathbf{m}$ and $\gcd(k,\mathbf{n})=1$.
  3. (iii)
    $\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b^{k}$ where $k\equiv 1\mod\mathbf{m}$ and $\gcd(k,\mathbf{n})=1$.
Proof.
The case $k=1$ is trivial, so we may assume $k>1$.
$(i)\Rightarrow(ii)$: If there exists an isomorphism $G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ with $k>1$ for some $\tau$ that commutes with $\sigma$ and some $\alpha\in S^{\times}$, then $k\equiv 1\mod\mathbf{m}$ and $\gcd(k,\mathbf{n})=1$ by Theorem 4.4 and we can construct a degree one isomorphism $G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b^{k})$ by Proposition 5.5 (a).
$(ii)\Rightarrow(i)$: If we have a degree one isomorphism $G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b^{k})$ such that $k\equiv 1\mod\mathbf{m}$, $\gcd(k,\mathbf{n})=1$ then $\tau(a)=N_{\mathbf{n}}^{\sigma}(\alpha)b^{k}$ implies that $\tau(a)=(N_{S/S_{0}}(\alpha))^{\mathbf{n}/\mathbf{m}}b^{k}$ holds, hence $G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism by Theorem 4.4.
$(ii)\iff(iii)$ This is Proposition 4.2. ∎
Corollary 5.10.
Suppose $K=\mathbb{F}_{p^{r}}$ and $\sigma(x)=x^{p^{s}}$ so that $\mathbf{m}=r/s$. Suppose that $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ is associative, that is, $a\in K_{0}^{\times}$ and $\mathbf{m}|\mathbf{n}$. Then the number of different $(\mathbf{n},\sigma)$-equivalence classes of skew $(\sigma,a)$-constacyclic codes equals the number of different $(\mathbf{n},\sigma)$-isometry classes of skew $(\sigma,a)$-constacyclic codes.
6. On non monomial homomorphisms
Having established the scarcity of monomial homomorphisms, we now show further that if $S$ is an integral domain (e.g., a field), then in a large number of cases every homomorphism between proper nonassociative algebras of the form $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is Hamming weight preserving — in fact, monomial of degree one, that is, is of the form $G_{\tau,\alpha}$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$ and nonzero $\alpha\in S$.
We begin by proving a generalization to the polynomial case of Lemma 4.3. We assume in this section that $S$ has no nonzero zero divisors, that is, it is an integral domain, such as a field.
Theorem 6.1.
Suppose $f(t)\in S[t;\sigma]$ is a monic polynomial of degree $\mathbf{n}$ and $b\in S^{\times}$. Suppose that $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a nonzero homomorphism such that $G(t)$ is not a monomial, whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$. Then $\mathbf{m}|\mathbf{n}$ and $G(t)=p(t^{\mathbf{m}})t$ for some $p\in S[x]$ of degree less than $\frac{\mathbf{n}}{\mathbf{m}}$.
Proof.
Write
$G(t)=\sum_{i=0}^{\mathbf{n}-1}c_{i}t^{i}$
for some $c_{i}\in S$. As in the proof of Lemma 4.3, which holds for any commutative unital ring $\S$, the condition $G(td)=G(\sigma(d)t)$ for all $d\in S$ implies that $c_{i}=0$ for all $i$ such that $i\not\equiv 1\mod\mathbf{m}$. Therefore $G(t)$ takes the form given in the statement. It remains to prove that $\mathbf{m}|\mathbf{n}$.
As compared with the monomial case, the proof now is more intricate because each term of $G(t)^{s}$ is a complex polynomial. In the following, we differentiate between the expression $(tp(t^{\mathbf{m}}))^{s}$ in $S[t;\sigma]$ and the value of $G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, that is, after reduction ${\rm mod}_{r}(t^{\mathbf{n}}-b)$ (whose well-definedness is the key question).
Let $k=\deg(G(t))>1$. Let $s>1$ be the least value for which $sk\geq\mathbf{n}$; since $k>1$ we have $s<\mathbf{n}$. Then since $G$ is homomorphism, $G(t)=tp(t^{\mathbf{m}})$ must be power associative for all powers less than $\mathbf{n}$ (since this is true in $S[t;\sigma]/S[t;\sigma]f(t)$) and we have $G(t^{s})=G(t)^{s}$. Furthermore, since $G(t)^{s-1}=(tp(t^{\mathbf{m}}))^{s-1}$ has degree less than $\mathbf{n}$, all monomials appearing with nonzero coefficient in $(tp(t^{\mathbf{m}}))^{s}\in S[t;\sigma]$ have degree less than $2\mathbf{n}$. Thus we may compute $G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ by first evaluating the expression $(tp(t^{\mathbf{m}}))^{s}$ in $S[t;\sigma]$ and then reducing modulo the relation $t^{\mathbf{n}+r}\equiv t^{r}b\equiv\sigma^{r}(b)t^{r}$.
Write $\mathcal{D}^{s}$ for the set of all $d\in\mathbb{N}$ that occur as the degree of some monomial appearing in the expansion of $(tp(t^{\mathbf{m}}))^{s}\in S[t;\sigma]$ and $\mathcal{B}^{s}$ for the set of all degrees appearing in $G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$. Then $\mathcal{B}^{s}\subset\{0,1,\cdots,\mathbf{n}-1\}$ and $\mathcal{D}^{s}\subset\mathcal{B}^{s}+(\mathbf{n}+\mathcal{B}^{s})$.
Since each monomial of $G(t)=tp(t^{\mathbf{m}})$ has degree of the form $1+\ell\mathbf{m}$ for some $\ell\geq 0$, it follows that for each $d\in\mathcal{D}^{s}$ we have $d\equiv s\mod\mathbf{m}$. For each such $d$ that satisfies $d\geq\mathbf{n}$, we obtain a monomial term $\lambda t^{d-\mathbf{n}}$ in the expression for $G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, for some $\lambda\in S$.
If $\lambda=0$, then some cancellation occurred, which necessarily implies that $d-\mathbf{n}=d^{\prime}$ for some $d^{\prime}\in\mathcal{D}^{s}$ satisfying $d^{\prime}<\mathbf{n}$. Subtracting $s$ from both sides yields $\mathbf{n}\equiv 0\mod\mathbf{m}$, or $\mathbf{m}|\mathbf{n}$.
If $\lambda\neq 0$, then we may write
$G(t)^{s}=\lambda t^{d-\mathbf{n}}+q(t)$
where $q(t)$ contains no monomials of degree $d-\mathbf{n}$. Let $c\in S$. Then the relation $t^{s}c=\sigma^{s}(c)t^{s}$ holds in the domain of $G$; since $G$ is a homomorphism, this implies $G(t)^{s}\tau(c)=\tau(\sigma^{s}(c))G(t)^{s}$. This scalar relation holds on each monomial term and so in particular one must have
$\lambda t^{d-\mathbf{n}}\tau(c)=\tau(\sigma^{s}(c))\lambda t^{d-\mathbf{n}}.$
The expression on the left evaluates to $\lambda\sigma^{d-\mathbf{n}}(\tau(c))t^{d-\mathbf{n}}$, whence, since $\lambda$ is not a zero divisor, we infer
$\sigma^{d-\mathbf{n}}(\tau(c))=\tau(\sigma^{s}(c))$
for all $c\in S$. Since $\tau$ commutes with $\sigma$, this implies $d-\mathbf{n}\equiv s\mod\mathbf{m}$, and so again we deduce $\mathbf{n}\equiv 0\mod\mathbf{m}$, or $\mathbf{m}|\mathbf{n}$, as required. ∎
Corollary 6.2.
If $a,b\in S^{\times}$ and $\mathbf{m}\nmid\mathbf{n}$, then any nonzero homomorphism $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$, must be monomial of degree one.
Proof.
The theorem shows that there are no polynomial homomorphisms in this case and Corollary 4.5 proves that under the same hypothesis plus any Hamming-weight-preserving homomorphism has degree one. ∎
We now have another more precise result about the nonexistence of nonmonomial homomorphisms, subject to a technical hypothesis ($\star$6.3).
Theorem 6.3.
Suppose $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a non monomial homomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$, as in Theorem 6.1. Let $k=\deg(G(t))$. Suppose further that
[…]
Then $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ must be associative, that is, $b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$.
For example, hypothesis ($\star$6.3) holds whenever $k<\mathbf{n}/2$.
Proof.
By Theorem 6.1, the degrees of each of the monomials appearing in $G(t)=tp(t^{\mathbf{m}})$ are congruent to $1$ mod $\mathbf{m}$, and $\mathbf{m}|\mathbf{n}$.
We can write $k=1+\ell\mathbf{m}=\deg(G(t))$. We have $1\leq k\leq\mathbf{n}-\mathbf{m}+1$. Since we assume $\mathbf{m}\geq 2$, we have $k\geq 3$ so $\mathbf{n}\geq 4$. Let $s>1$ be the least value for which $sk\geq\mathbf{n}$; considering $k=1+\mathbf{m}$ gives the upper bound $s<\frac{\mathbf{n}}{\mathbf{m}+1}+1\leq\frac{\mathbf{n}}{3}+1$. Since $s$ is an integer this implies for all $\mathbf{n}\geq 4$ that $s\leq\mathbf{n}-2$. Therefore $s+1<\mathbf{n}$ and since $G$ is a homomorphism we have $G(t^{s+1})=G(t)^{s+1}$. We now compute $G(t)^{s+1}$ in two ways: as $G(t)^{s}G(t)$ and as $G(t)G(t)^{s}$.
We first set some notation. Let $\mathcal{D}^{s+1}$ be the set of all degrees of monomials appearing in $H(t)=(t(p(t^{\mathbf{m}}))^{s+1}\in S[t;\sigma]$. Then the maximal element of $\mathcal{D}^{s+1}$ is $d_{max}=(s+1)k$. Write
$G(t)=\sum_{i=0}^{\ell-1}c_{i}t^{1+i\mathbf{m}}$
for some $c_{i}\in S$. For any $d$, let
$I_{d}=\{\gamma=(\gamma_{1},\gamma_{2},\cdots,\gamma_{s+1})\mid c_{\gamma_{j}}\neq 0,\sum_{j}(1+\gamma_{j}\mathbf{m})=d\}.$
These index all the different products of terms of $G(t)$ that give rise to a monomial of degree $d$ in the product $H(t)\in S[t;\sigma]$. Specifically, given $\gamma\in I_{d}$, the corresponding (noncommutative!) product is
$c_{\gamma}t^{d}=\prod_{j=1}^{s+1}(c_{\gamma_{j}}t^{1+\gamma_{j}\mathbf{m}})=\prod_{j=1}^{s+1}\sigma^{j-1}(c_{\gamma_{j}})t^{d}$
where this is well-defined in $S[t;\sigma]$. If $d<\mathbf{n}$, then $c_{\gamma}$ is also the coefficient corresponding to this monomial product in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, since products with total degree less than $\mathbf{n}$ are associative. If $d>\mathbf{n}$, however, then the value of the coefficient will a priori depend on the order of associators.
By ($\star$6.3), we have $(s+1)k<2\mathbf{n}$. Then $d^{\prime}=d_{max}-\mathbf{n}=(s+1)k-\mathbf{n}$ is the degree of a monomial appearing in $G(t)^{s+1}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ and we wish to compare its coefficients as computed in the two distinct ways ($G(t)G(t)^{s}$ and $G(t)^{s}G(t)$). The coefficient of $t^{d^{\prime}}$ in $G(t)^{s+1}$ in either case is the sum of all monomials in $I_{d^{\prime}}$, plus additional terms that correspond to products in $I_{d_{max}}$ that were reduced by the relation in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ to terms of degree $d^{\prime}$. By construction, however, $I_{d_{max}}$ is a singleton, being composed only of the products of the top degree terms:
$I_{d_{max}}=\{(\ell,\ell,\cdots,\ell)\}=\{\gamma_{max}\}.$
We evaluate the corresponding term in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ in two ways as follows. First note that
$(c_{\ell}t^{k})^{s}=N_{s}^{\sigma}(c_{\ell})t^{sk}=N_{s}^{\sigma}(c_{\ell})t^{sk-\mathbf{n}}b=N_{s}^{\sigma}(c_{\ell})\sigma^{s}(b)t^{sk-\mathbf{n}}$
where we have used right division by $t^{\mathbf{n}}-b$ in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ and that $sk-\mathbf{n}\equiv s\mod\mathbf{m}$. Thus the monomial corresponding to $\gamma_{max}$ in the product $G(t)^{s}G(t)$ is
$c_{left}t^{d^{\prime}}=(c_{\ell}t^{k})^{s}(c_{\ell}t^{k})=N_{s}^{\sigma}(c_{\ell})\sigma^{s}(b)t^{sk-\mathbf{n}}c_{\ell}t^{k}=\sigma^{s}(b)N_{s+1}^{\sigma}(c_{\ell})t^{(s+1)k-\mathbf{n}}$
whereas the monomial corresponding to $\gamma_{max}$ in the product $G(t)G(t)^{s}$ is
$c_{right}t^{d^{\prime}}=(c_{\ell}t^{k})(c_{\ell}t^{k})^{s}=c_{\ell}t^{k}N_{s}^{\sigma}(c_{\ell})\sigma^{s}(b)t^{sk-\mathbf{n}}=N_{s+1}^{\sigma}(c_{\ell})\sigma^{s+1}(b))t^{(s+1)k-\mathbf{n}}.$
Therefore, since $G(t)^{s}G(t)=G(t)G(t)^{s}$, the coefficients of $t^{d^{\prime}}$ in each of these expressions must be equal, yielding
$\sum_{\gamma\in I_{d^{\prime}}}c_{\gamma}+c_{left}=\sum_{\gamma\in I_{d^{\prime}}}c_{\gamma}+c_{right}.$
As the sums are equal this reduces to the requirement that $c_{left}=c_{right}$, which implies by cancellability of $c_{\ell}\neq 0$ that $\sigma^{s}(b)=\sigma^{s+1}(b)$ or $b\in S_{0}$. ∎
Corollary 6.4.
Suppose $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a non monomial isomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$, and whose degree $k$ satisfies ($\star$6.3), as in Theorem 6.1. Then $a,b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$.
On the other hand, in the associative case it is not difficult to produce a homomorphism $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ for which $G(t)$ is a non monomial polynomial.
Example 6.5.
The minimal example is $\mathbf{m}=2,\mathbf{n}=4$ and $G(t)=t+t^{3}$, with $\tau=\operatorname{id}$. We choose $K=\mathbb{F}_{25}$ and $K_{0}=\mathbb{F}_{5}$ with $\sigma(x)=x^{5}$. Then $\mathbf{m}|\mathbf{n}$ and $a=b=4\in K_{0}$, so $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ is associative. Since in addition $G(t)^{2}=(t+t^{3})^{2}=t^{2}+2t^{4}+4t^{2}=3$, we have $G(t)^{4}=4$ so $G$ is a well-defined (non-surjective!) polynomial morphism
$G:K[t;\sigma]/K[t;\sigma](t^{4}-4)\to K[t;\sigma]/K[t;\sigma](t^{4}-4).$
The hypothesis ($\star$6.3) is a minor technical condition that should not be necessary. We present work towards proving the following conjecture in [NP].
Conjecture 6.6.
Suppose $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a non monomial isomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$. Then $a,b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$, that is, $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ must be associative.
As mentioned, having such a complete understanding of all homomorphisms between Petit algebras has significant impact to the classification of various associated error-correcting codes. In addition to the conjecture above, we identify that it is possible to extend the results of our paper to the case of noncommutative rings. It would also be interesting to see how our methods can be applied when we look at homomorphisms between general Petit algebras.
Acknowledgments
This paper was written during the second author’s stay as a Simons Professor in Residence at the University of Ottawa. She gratefully acknowledges the support of CRM and the Simons Foundation.